# Difficulty-Aware Agent Orchestration in LLM-Powered Workflows

### v2

@@ Abstract
Large Language Model (LLM)-based agentic systems have shown strong capabilities across various tasks. However, existing multi-agent frameworks often rely on static or task-level workflows, which either over-process simple queries or underperform on complex ones, while also neglecting the efficiency-performance trade-offs across heterogeneous LLMs. To address these limitations, we propose Difficulty-Aware Agentic Orchestration (DAAO), a dynamic framework that adapts workflow depth, operator selection, and LLM assignment based on the difficulty of each input query. DAAO comprises three interdependent modules: a variational autoencoder (VAE) for difficulty estimation, a modular operator allocator, and a cost- and performance-aware LLM router. By leveraging heterogeneous LLMs and dynamically tailoring workflows, DAAO enables fine-grained, query-specific reasoning strategies. DAAO outperforms prior multi-agent systems in both accuracy and inference efficiency across six benchmarks. We will release our code and implementation details upon publication.
Figure 1: The overall framework of our proposed DAAO.
Introduction
@@
Large Language Model (LLM)-based agents (Richards and et al. 2023; Nakajima 2023; Reworkd 2023) have exhibited remarkable capabilities across a wide spectrum of tasks, including question answering (Zhu et al. 2024), data analysis (Hong et al. 2024; Li et al. 2024), decision-making (Song et al. 2023), code generation (Shinn, Labash, and Gopinath 2023) and web navigation (Deng et al. 2024). Building upon the success of single agents, recent advancements reveal that organizing multiple LLM-based agents into structured agentic workflows can significantly enhance task performance. In such workflows, agents can interact either cooperatively (Zhuge et al. 2024) or competitively (Zhao et al. 2023) depending on the task context. These multi-agent systems can overcome the cognitive and functional limitations of individual models (Du et al. 2023; Liang et al. 2023; Wang et al. 2023b; Jiang, Ren, and Lin 2023; Wu et al. 2023; Zhang et al. 2024a), thereby exhibiting collective intelligence similar to human collaboration in a society of agents.
@@
In recent years, the research community has focused on automating multi-agent system design. For instance, DsPy (Khattab et al. 2023) and EvoPrompting (Guo et al. 2023) automate prompt optimization, GPTSwarm (Zhuge et al. 2024) optimizing inter-agent communication, and EvoAgent (Yuan et al. 2024) self-evolving agent profiling. However, these systems are often constrained by limited search spaces and rigid representation paradigms, resulting in marginal performance gains and limited adaptability to diverse task requirements. Subsequently, ADAS (Hu, Lu, and Clune 2024) and AFlow (Zhang et al. 2024b) employ code as representation for workflow, facilitating robust and flexible workflow searches through different paradigms, with ADAS utilizing heuristic search and AFlow adopting Monte Carlo tree search. MaAS (Zhang et al. 2025) proposes an agentic supernet to generate a query-specific multi-agent system for each user query.
@@
Despite their success, existing automation pipelines often lack complex adaptivity and LLM heterogeneity. Task-level workflows (Hu, Lu, and Clune 2024; Zhang et al. 2024b) are typically built as uniform multi-agent systems for entire task categories, achieving strong metrics like accuracy and pass@k but relying on heavy pipelines with excessive LLM calls and tool usage. This design over-processes simple queries, wasting resources and overlooking factors like token cost and latency. Query-level workflows (Zhang et al. 2025) introduce input-specific adaptation, but their granularity is often insufficient, leading to suboptimal or oversimplified workflows for difficult inputs. These limitations motivate a difficulty-adaptive framework that dynamically balances complexity and cost. Furthermore, most workflows (Hu, Lu, and Clune 2024; Zhang et al. 2024b, 2025) rely on a single large LLM (e.g., GPT-4o (OpenAI 2024)), ignoring recent findings (Ye et al. 2025; Chen, Zaharia, and Zou 2023; Hu et al. 2024) that different LLMs exhibit complementary capabilities. Smaller models can outperform larger ones on specific tasks while significantly reducing cost. Thus, there is increasing support for agentic workflows that integrate diverse LLMs of varying sizes, leveraging their specialization to improve performance and efficiency.
To address the challenges of adapting reasoning strategies to queries with varying difficulty and domain characteristics, we propose Difficulty-Aware Agentic Orchestration (DAAO). DAAO is a dynamic framework that estimates the difficulty of each incoming query and composes optimized workflows by selecting suitable agentic operators—such as Chain-of-Thought (CoT) (Wei et al. 2022), Multi-Agent Debate (Du et al. 2023), and ReAct (Yao et al. 2023)—and assigning them to appropriate LLMs. By tailoring both workflow depth and model assignment based on query characteristics, DAAO balances reasoning effectiveness and computational cost. Furthermore, the system refines its predictions over time through feedback, enabling continual adaptation.
Technically, DAAO decomposes the orchestration process into three interdependent modules. It begins with a query difficulty estimator, implemented as a variational autoencoder (VAE) (Kingma and Welling 2014), which encodes each input query into a latent representation reflecting its difficulty. This difficulty signal then guides the agentic operator allocator, which determines the depth of the workflow and selects suitable reasoning operators based on both the query and its inferred complexity. Finally, the LLM router assigns each selected operator to a language model that best fits its role, enabling heterogeneous and cost-effective reasoning. Together, these modules form a cohesive and adaptive orchestration framework that supports flexible, query-specific agentic workflows.
We conduct comprehensive evaluations on six widely adopted benchmarks. Empirical results reveal that our DAAO achieves both state-of-the-art performance and remarkable cost-reduction. For instance, our DAAO surpasses state-of-the-art automated multi-agent systems by $11.21\%$ in accuracy, while requiring only $64\%$ of their inference costs, owing to its difficulty-aware routing and adaptive workflow construction.
@@
  • Query-Level Difficulty Estimation for Workflow Adaptation: We propose a difficulty-aware orchestration framework that estimates the complexity of each input query using a variational autoencoder (VAE). This latent difficulty signal informs both the depth and structure of the workflow, enabling fine-grained, query-specific adaptation beyond static or task-level designs.
  • Modular Agentic Workflow Construction with Heterogeneous Operators and LLMs: We develop a modular pipeline that selects agentic operators based on query difficulty and context, and assigns each to the most suitable LLM. This fine-grained routing leverages the complementary strengths of diverse models to enhance reasoning performance while reducing cost.
  • Efficient and Scalable Execution with Empirical Gains: Our method achieves state-of-the-art performance on six benchmarks, outperforming prior multi-agent systems by up to 11.21% in accuracy while using only 64% of the inference cost. The effectiveness of difficulty-guided orchestration is further validated through comprehensive ablation and cost-efficiency analysis.
Figure 2: Case study and visualization for MaAS (Zhang et al. 2025) and Ours. Queries are from MATH, HumanEval and GAIA benchmarks. Left: our method solves the query while MaAS fails. Middle: our workflow is simpler for easy queries. Right: our method leverages heterogeneous LLMs within the same workflow structure.
Related Work
@@
Automated Agentic Workflows. The development of agentic workflows has evolved from manual configurations to automated systems, with the latter offering improved adaptability and task performance. Early approaches to automation focus on optimizing prompt structures and inter-agent communication protocols (Khattab et al. 2023; Guo et al. 2023; Zhuge et al. 2024; Yuan et al. 2024), thereby enhancing the robustness of workflows across a variety of tasks. More recent systems, such as ADAS (Hu, Lu, and Clune 2024) and AFlow (Zhang et al. 2024b), leverage code-based representations to enable real-time structural adaptation and communication strategy refinement based on environmental feedback. MaAS (Zhang et al. 2025) further introduces query-specific multi-agent composition using a supernet-like architecture. Despite these advances, most existing workflows remain LLM-homogeneous, typically instantiating all agents with a single high-capacity model (e.g., GPT-4o (OpenAI 2024)), which limits both efficiency and the ability to specialize across sub-tasks. Additionally, the adaptivity of these systems is often coarse-grained, leading to under- or over-processing depending on query complexity.
LLM Heterogeneity and Difficulty-Aware Reasoning. Emerging research increasingly highlights the complementary strengths of LLMs with different capacities. Works such as X-MAST (Ye et al. 2025), FrugalGPT (Chen, Zaharia, and Zou 2023), and RouterBench (Hu et al. 2024) demonstrate that smaller LLMs can outperform larger models in specific domains while incurring lower inference costs. These findings suggest that heterogeneous model ensembles, if properly orchestrated, can achieve better cost-performance trade-offs than homogeneous systems (Gong et al. 2023). Concurrently, there is growing interest in difficulty-aware reasoning, where model behavior is conditioned on the estimated hardness of each query (Li et al. 2023; Zhang et al. 2023). However, most existing systems do not incorporate fine-grained difficulty estimation into workflow generation (Lin et al. 2023). Our work builds upon these insights by introducing a unified framework that integrates both difficulty-aware reasoning and LLM heterogeneity for efficient and adaptive agentic workflow construction.
Preliminary
In this section, we formally define the search space of DAAO and its associated optimization objective.
Search Space.
We define the basic unit of DAAO’s search space, referred to as an agentic operator. An agentic operator $\mathcal{O}$ is a composite module consisting of a language model and an associated collaborative protocol:
$\begin{gathered}\mathcal{O}_{j}^{i}=\{\mathcal{M},\mathcal{P}\},\\ \mathcal{M}\in\mathbb{M},\quad\mathcal{P}\in\mathbb{P},\end{gathered}$ (1)
@@
where $\mathcal{M}$ denotes a selected LLM backbone from the set of available models $\mathbb{M}$, and $\mathcal{P}$ specifies a collaboration protocol from the protocol set $\mathbb{P}$. For example, $\mathcal{O}_{\text{CoT}}^{q}=\{\text{Qwen2-70B},\text{Chain-of-Thought}\}$ denotes using Qwen2-70B with a step-by-step reasoning strategy, whereas $\mathcal{O}_{\text{debate}}^{g}=\{\text{GPT-4o-mini},\text{Multi-agent debate}\}$ configures multiple GPT-4o-mini instances to engage in turn-based argumentation. The set of feasible operators is denoted by $\mathbb{O}$.
We define an agentic workflow as a directed acyclic graph (DAG):
$\mathcal{G}=\{\mathcal{V},\mathcal{E}\},\quad\mathcal{V}\subset\mathbb{O},\quad\mathcal{E}\subset\mathcal{V}\times\mathcal{V},$ (2)
where $\mathcal{V}$ denotes the set of selected operators and $\mathcal{E}$ encodes their directed connectivity. The DAG structure enforces hierarchical operator execution.
Definition of DAAO.
DAAO is denoted as $\mathcal{A}=\{\boldsymbol{\pi},\mathbb{O}\}=\{\{\pi_{\ell}(\mathcal{O})\}_{\mathcal{O}\in\mathbb{O}}\}_{\ell=1}^{L}$, where
$\pi_{\ell}(\mathcal{O})$ $=p(\mathcal{O}\mid\mathcal{A}_{1:\ell-1}),\quad\mathcal{O}\in\mathbb{O},$ (3)
$\mathcal{A}_{1:\ell-1}$ $=\left\{\{\pi_{k}(\mathcal{O})\}_{\mathcal{O}\in\mathbb{O}}\right\}_{k=1}^{\ell-1},$
and $\pi_{\ell}(\mathcal{O})$ denotes the probability of selecting operator $\mathcal{O}$ at layer $\ell$, conditioned on the prior layers $\mathcal{A}_{1:\ell-1}$. This defines a sequential decision process for operator selection.
The induced supernet defines a joint distribution over all multi-layer agentic workflow configurations:
$p(\mathcal{G})=\prod_{\ell=1}^{L}\prod_{\mathcal{O}\in\mathbb{O}}\pi_{\ell}(\mathcal{O})^{\mathbb{I}_{\mathcal{O}\in\mathcal{V}_{\ell}}},$ (4)
where $\mathbb{I}_{\mathcal{O}\in\mathcal{V}_{\ell}}$ is an indicator function denoting the inclusion of operator $\mathcal{O}$ at layer $\ell$ of the workflow.
@@
Given a benchmark dataset $\mathcal{D}$ containing queries $\mathcal{Q}$ and their corresponding oracle answers $a$, the objective of DAAO is to learn a query-conditioned policy that balances task utility and inference cost. The objective is formally defined as:
@@
$\underset{\mathbb{P}(\mathcal{G}\mid\mathcal{Q})}{\max}\mathbb{E}_{\begin{subarray}{c}(\mathcal{Q},a)\sim\mathcal{D}\\ \mathcal{G}\sim\mathbb{P}(\mathcal{G}\mid\mathcal{Q})\end{subarray}}\left[U(\mathcal{G};\mathcal{Q},a)-\lambda\cdot C(\mathcal{G};\mathcal{Q})\right],\quad\text{s.t.}\mathcal{G}\subset\mathcal{A},$ (5)

### v3

# Difficulty-Aware Agentic Orchestration for Query-Specific Multi-Agent Workflows

@@ Abstract.
Large Language Model (LLM)-based agentic systems have shown strong capabilities across various tasks. However, existing multi-agent frameworks often rely on static or task-level workflows, which either over-process simple queries or underperform on complex ones, while also neglecting the efficiency-performance trade-offs across heterogeneous LLMs. To address these limitations, we propose Difficulty-Aware Agentic Orchestration (DAAO), which can dynamically generate query-specific multi-agent workflows guided by predicted query difficulty. DAAO comprises three interdependent modules: a variational autoencoder (VAE) for difficulty estimation, a modular operator allocator, and a cost- and performance-aware LLM router. A self-adjusting policy updates difficulty estimates based on workflow success, enabling simpler workflows for easy queries and more complex strategies for harder ones. Experiments on six benchmarks demonstrate that DAAO surpasses prior multi-agent systems in both accuracy and inference efficiency, validating its effectiveness for adaptive, difficulty-aware reasoning.
Multi-agent system, Difficulty-Aware, Adaptive workflows
$c$s: Computing methodologies Cooperation and coordinationccs: Computing methodologies Multi-agent systems
Figure 1. The overall framework of our proposed DAAO.
1. Introduction
@@
Large Language Model (LLM)-based agents (Richards and et al., 2023; Nakajima, 2023; Reworkd, 2023) have exhibited remarkable capabilities across a wide spectrum of tasks, including question answering (Zhu et al., 2024), data analysis (Hong et al., 2024; Li et al., 2024), decision-making (Song et al., 2023), code generation (Shinn et al., 2023) and web navigation (Deng et al., 2024). Building upon the success of single agents, recent advancements reveal that organizing multiple LLM-based agents into structured agentic workflows can significantly enhance task performance. In such workflows, agents can interact either cooperatively (Zhuge et al., 2024) or competitively (Zhao et al., 2023) depending on the task context. These multi-agent systems can overcome the cognitive and functional limitations of individual models (Du et al., 2023; Liang et al., 2023; Wang et al., 2023a; Jiang et al., 2023; Wu et al., 2023; Zhang et al., 2024b), thereby exhibiting collective intelligence similar to human collaboration in a society of agents.
@@
In recent years, the research community has focused on automating multi-agent system design. For instance, DsPy (Khattab et al., 2023) and EvoPrompting (Guo et al., 2023) automate prompt optimization, GPTSwarm (Zhuge et al., 2024) optimizing inter-agent communication, and EvoAgent (Yuan et al., 2024) self-evolving agent profiling. However, these systems are often constrained by limited search spaces and rigid representation paradigms, resulting in marginal performance gains and limited adaptability to diverse task requirements. Subsequently, ADAS (Hu et al., 2024) and AFlow (Zhang et al., 2024a) employ code as representation for workflow, facilitating robust and flexible workflow searches through different paradigms, with ADAS utilizing heuristic search and AFlow adopting Monte Carlo tree search. MaAS (Zhang et al., 2025) proposes an agentic supernet to generate a query-specific multi-agent system for each user query.
@@
How to dynamically generate a workflow given a query remains a key challenge in current research. Task-level workflows (Hu et al., 2024; Zhang et al., 2024a) are typically built as uniform multi-agent systems for entire task categories, achieving strong metrics like accuracy and pass@k but relying on heavy pipelines with excessive LLM calls and tool usage. This design over-processes simple queries, wasting resources and overlooking factors like token cost and latency. Query-level workflows (Zhang et al., 2025) introduce input-specific adaptation, but their granularity is often insufficient, leading to suboptimal or oversimplified workflows for difficult inputs. For instance, when a user requests a travel guide for a specific location, a workflow that only retrieves and summarizes information often falls short of meeting the user’s needs. These limitations motivate a difficulty-adaptive framework that dynamically balances complexity and cost.
To address the above challenges, we propose Difficulty-Aware Agentic Orchestration (DAAO), which can intelligently generate workflows according to the characteristics of each query. DAAO has three core capabilities: (1) Learning to capture the difficulty of each query from posterior knowledge, without relying on manual labels; (2) Dynamically creating workflows that match the predicted difficulty of each query; (3) Assigning LLMs to workflow components to maximize the reasoning ability of the Multi-agent system.
Technically, we define query difficulty as a learnable policy. Unlike previous methods, LLMs have little knowledge about workflow generation, and manually creating query-workflow pairs requires much human effort, which goes against automatic workflow generation. To address this, We use a reward-like mechanism to update the policy. When a workflow successfully solves a query, we slightly lower its predicted difficulty, allowing future workflows to be simpler. If a workflow fails, we increase the predicted difficulty to encourage more complex and capable workflows. In addition, we model multi-agent workflow generation on the agentic net, a probabilistic, continuous agentic architecture distribution that encompasses a vast number of possible multi-agent candidates. To enhance efficiency and adaptability, we incorporate a cost- and performance-aware LLM router that dynamically assigns heterogeneous models to different operators according to query difficulty and resource constraints. During training, a controller network samples multi-agent architectures conditioned on the input query and updates its policy through feedback signals. During inference, for different queries, DAAO samples a suitable multi-agent system delivering satisfactory resolution and appropriate inference resources.
@@
  • Query-Level Difficulty Estimation for Workflow Adaptation. We propose generating adaptive workflow strategies guided by query difficulty. The framework learns to represent the latent difficulty space of queries, leveraging workflow feedback to adjust the strategies.
  • Dynamic Workflow Generation. We propose DAAO, a difficulty-aware framework that dynamically generates workflows and realizes LLM heterogeneity based on query difficulty, domain, and features, while achieving performance-cost balance.
  • Experimental Validation. We conduct comprehensive evaluations on six widely adopted benchmarks, covering diverse use cases in code generation (HumanEval, MBPP), mathematical reasoning (GSM8K, MATH), knowledge and reasoning understanding (MMLU) and diverse tool usage (GAIA). Empirical results demonstrate that DAAO is (1) highly performing, surpassing existing automated orchestration methods by $3.5\%\sim 15.2\%$ and recent LLM routing methods by $3.2\%\sim 10.2\%$; (2) economical, outperforming the SOTA baseline MasRouter on the MATH benchmark with 65% of the training cost and 41% of the inference cost; (3) inductive, demonstrating strong generalization to unseen LLM backbones and transferability across diverse datasets.
2. Related Work
Automated Agentic Workflows.
@@ 2. Related Work
The development of agentic workflows has evolved from manual configurations to automated systems, with the latter offering improved adaptability and task performance. Early approaches to automation focus on optimizing prompt structures and inter-agent communication protocols (Khattab et al., 2023; Guo et al., 2023; Zhuge et al., 2024; Yuan et al., 2024), thereby enhancing the robustness of workflows across a variety of tasks. More recent systems, such as ADAS (Hu et al., 2024) and AFlow (Zhang et al., 2024a), leverage code-based representations to enable real-time structural adaptation and communication strategy refinement based on environmental feedback. MaAS (Zhang et al., 2025) further introduces query-specific multi-agent composition using a supernet-like architecture. Despite these advances, current frameworks still face two major limitations. First, most multi-agent frameworks remain LLM-homogeneous, relying on a single backbone model (e.g., GPT-4o-mini) for all agents and thus missing the benefits of heterogeneous collaboration across models. Second, they lack complexity diversity: most systems adopt uniformly complex workflows optimized for accuracy, ignoring that real-world queries vary widely in difficulty.
Difficulty-Aware Reasoning.
Recent advances in the reasoning domain of large language models increasingly emphasize difficulty-aware mechanisms to address the limitations of uniform reward signals across heterogeneous tasks, particularly in mathematical reasoning where problem complexity varies widely (Li et al., 2025; Zhang and Zuo, 2025; Ji et al., 2025; Tong et al., 2024). These methods dynamically adjust learning objectives based on task difficulty estimates, prioritizing deeper exploration for challenging problems while promoting efficiency on simpler ones. However, the model itself lacks intrinsic knowledge of the difficulty of multi-agent workflows, making it unable to align the difficulty of queries with that of multi-agent workflows. To address this, we propose incorporating difficulty awareness into automatic workflow generation, resolving the perception of workflow difficulty for queries and enhancing workflow adaptability.
3. Methodology
Overview.
Figure 1 illustrates our Difficulty-Aware Agentic Orchestration (DAAO), which generates query-specific agentic workflows across domains and difficulty levels. Our key contribution is a standalone query difficulty estimator $N_{\theta_{d}}$ that provides an explicit, calibrated difficulty signal for each input. Unlike prior controller networks (Zoph and Le, 2017) that score architectures without reliably assessing the query’s difficulty, $N_{\theta_{d}}$—instantiated as a variational autoencoder (VAE) (Kingma and Welling, 2014) with a learned difficulty head—encodes the query into a latent representation $z$ and outputs a scalar difficulty $d\in(0,1)$. This difficulty estimate conditions (i) a layered operator allocator $N_{\theta_{o}}$ that selects an appropriate subset of agentic operators and workflow depth, and (ii) a cost-aware LLM router $N_{\theta_{m}}$ that assigns backbone models by balancing reasoning needs with computational budget. The three modules together yield a customized multi-stage workflow per query. After execution, we evaluate the output quality and use the success signal to update $N_{\theta_{d}}$ and refine $N_{\theta_{o}}$/ $N_{\theta_{m}}$, enabling continual improvement while keeping difficulty estimation central to workflow construction.
3.1. Preliminary
This section formalizes the search space for difficulty-aware agentic workflow generation and the cost–utility objective optimized by our policy.
Agentic operator and workflow
Let $\mathbb{M}$ be the set of available large language models (LLMs) and $\mathbb{S}$ the set of collaboration protocols (e.g., Chain of Thought, Debate, Ensemble). The catalog of feasible operators is the subset $\mathbb{O}\subseteq\mathbb{M}\times\mathbb{S}$. An agentic operator is a pair of one model and one protocol:
(1) $O\;=\;\{M,S\},\qquad M\in\mathbb{M},\ \ S\in\mathbb{S},\ \ O\in\mathbb{O}.$
@@ 3.1. Preliminary
For example, $\{\text{Qwen2-72B},\text{Chain-of-Thought}\}$ denotes step-by-step reasoning on Qwen2-70B, whereas $\{\text{GPT-4o-mini},\text{Debate}\}$ configures turn-based multi-agent debate.
An agentic workflow can be described as a Directed Acyclic Graph (DAG):
(2) $G=\{\mathcal{V},\mathcal{E}\},\quad\mathcal{V}\subseteq\mathbb{O},\ \ \mathcal{E}\subseteq\mathcal{V}\times\mathcal{V},$
where $\mathcal{V}$ collects instantiated operators (nodes) and $\mathcal{E}$ encodes directed dependencies (edges). We endow each workflow with a topological layering $\mathcal{V}=\bigcup_{l=1}^{L}\mathcal{V}_{l}$ and restrict edges to go from earlier to later layers; i.e., if $u\in\mathcal{V}_{l}$ and $v\in\mathcal{V}_{l^{\prime}}$, then $l^{\prime}>l$.
Layered Policy
We define the layered selection policy of DAAO over the operator library $\mathbb{O}$ as
(3) $\mathcal{A}\;=\;\big\{\{\pi_{l}(O)\}_{O\in\mathbb{O}}\big\}_{l=1}^{L},\qquad\pi_{l}(O)\;=\;\mathbb{P}\!\big(O\,\big|\,\mathcal{Q},z\big),\ \ O\in\mathbb{O},$
where $z$ is the latent embedding produced by our difficulty estimator encoded from the input query $\mathcal{Q}$ and adaptively adjusted across layers l=1:Ll=1{:}L. In other words, $z=f_{\phi}(\mathcal{Q},\mathcal{H}_{1:l-1})$ is a learned state summary that encodes the query and the preceding layer-wise history $\mathcal{H}_{1:l-1}=\{\mathcal{A}_{k}\}_{k=1}^{l-1}$ of active operator sets ($\mathcal{A}_{k}\subseteq\mathbb{O}$). The policy induces a joint distribution over multi-layer operator configurations:
(4) $\mathbb{P}(G\mid\mathcal{Q})\;=\;\prod_{l=1}^{L}\prod_{O\in\mathbb{O}}\big(\pi_{l}(O\mid\mathcal{Q},z)\big)^{\mathbb{I}[\,O\in\mathcal{A}_{l}\,]},$
where $G=\{\mathcal{V},\mathcal{E}\}$ is the DAG in Eq. (2), and the node set is the union of layerwise active sets, respecting the topological layering.
@@
Given a benchmark dataset $\mathcal{D}$ containing queries $\mathcal{Q}$ and their oracle answers $a$, the objective of DAAO is to learn a query-conditioned policy that balances task utility and inference cost:
@@
(5) $\underset{\mathbb{P}(G\mid\mathcal{Q})}{\max}\ \mathbb{E}_{\begin{subarray}{c}(\mathcal{Q},a)\sim\mathcal{D}\\[2.0pt] G\sim\mathbb{P}(G\mid\mathcal{Q})\end{subarray}}\Big[\,U(G;\mathcal{Q},a)\;-\;\lambda\,C(G;\mathcal{Q})\,\Big],\quad\text{s.t.}\ \lambda\geq 0,$